Please help me solve me with this problem. A school had 900 students at the start of a school year. If there is an enrollment increase of 7% by the beginning of the next year, what is the new enrollment? : Please help me solve me with this problem. A school had 900 students at the start of a school year. If there is an enrollment increase of 7% by the beginning of the next year, what is the new enrollment?
Answer by stanbon(19705)

A school had 900 students at the start of a school year. If there is an enrollment increase of 7% by the beginning of the next year, what is the new enrollment?
---------------
New enrollment = 900 + 0.07*900 = 1.07*900 = 963 students
==================================================================
Cheers,
Stan H.